Incident X-rays have a wavelength 0.2169 nm and ar
Answer #1

Momentum of photon is given as

p=\frac{h}{\lambda}

A.

There fore for incident photon

p=\frac{6.626*10^{-34}}{0.2169*10^{-9}}=3.05*10^{-24}. N-s

B.

Wavelength of scattered photon

\lambda _{s} = \lambda_{I} +\frac{h(1-cos\theta)}{m_{e}c^2}

\lambda_{s}= 0.2169*10^{-9} + 1.35*10^{-20} m

Thus the momentum will be almost equal to the momentum of incident photon

= 3.05×10-24 N-s